Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook page YouTube channel

BlueSky profile YouTube channel
P68-027

P68-027: The 3DUP activity of AGB stars in fluorine nucleosynthesis studies

P.I.: Saberi

This is a Monitoring program using FIES.

Instrument setup

  • FIES

  • Med-res fiber F3

Afternoon instructions

  • Run afternoon calibrations: fies-calibs -fib 3

Observing instructions

  • On the Queue page, make sure the OB-generated observing script is written to disk, or else 'compile'

  • Go to directory: cd ~/obs/scripts/68-027/ and do ls -lrt

  • Start the requested OB generated script


Program Tracking
Proposal 68-027
Semester P68: 20231001 - 20240401
Principal Investigator Maryam Saberi
Programme The 3DUP activity of AGB stars in fluorine nucleosynthesis studies
Instrument FIES
Time Allocation 11 targets
Total Nights 5.5h
Type Mon
Notes Fixed dates to be scheduled
Total Time Spent 03:09 (hh:mm)
Proposal Date Time Type Targets Files Comments Status
68-027 2023-10-07 01:16 Mon AQSgr, RTCap, VAql FIGj070094 - FIGj070096 Observed all three targets with clear skies or light haze. Dusty conditions (> 35 micro grams per m3) all around and Dimm seeing below 1.4 arcsec in general during the observations. I had slight difficulties during the acquisition of target rt_cap as the target seemed to be quite off-center in the frames, but after aborting the script twice, and getting the same result I decided to proceed the third time. Observed
68-027 2023-10-11 01:33 Mon rv_aqr, rv_cyg FIGj110095 - FIGj110096 , FIGj110097 - FIGj110098 The first attempt to get rv_aqr had to be aborted due to clouds. Later, both targets were observed in good conditions. Observed
68-027 2023-10-14 00:20 Mon RScl FIGj140107 - FIGj140107 Observed
68-027 2024-01-01 00:00 Mon All observations were missed due to telescope being closed because of high humidity. Failed
68-027 2024-02-02 00:00 Mon Both targets were attempted but could not be observed due to ongoing issue with stancam (#FR 4410). The fiber guiding is so out of focus that the fiber-head was not visible. Failed

 

The above information is collected from the ToO/Mon reports that have been created by the observer.

Should any of the observations have been carried out using the OB system, you will find additional information HERE

Back to top Last modified: September 27 2023